# Bones of Contention: Exploring Query-Efficient Attacks Against Skeleton Recognition Systems

### v1

@@ IV-B Perturbation Optimization
Input: Black-box model $f$, surrogate model $f_{s}$, key joint number $N_{k}$, input motion $\mathbf{x}$, original class $y$, intra constraint threshold $\varepsilon_{b}$, tolerance $\gamma$.
Output: Adversarial motion $\mathbf{x}_{adv}$.
1 $M\leftarrow\operatorname{\texttt{key\_joint}}\left(f_{s},N_{k},\mathbf{x},y\right)$;
2 $\mathbf{d}_{best}\leftarrow M\odot\left[1,1,...,1\right]^{T}$, $r_{best}\leftarrow\infty$, $s\leftarrow 0$;
3 while not reaching the query limit do
4       $\mathbf{d}_{cur}\leftarrow\mathbf{d}_{best}$;
5       Split $\mathbf{d}_{cur}$ into $2^{s}$ blocks with index $I_{0},I_{1},...,I_{2^{s}}$;
6       for $k\leftarrow 0$ to $2^{s}$ do
7             $\mathbf{d}_{cur}\left[I_{k}\right]\leftarrow-\mathbf{d}_{cur}\left[I_{k}\right]$;
8             $\mathbf{d}_{nor}\leftarrow\mathbf{d}_{cur}/{\left\|\mathbf{d}_{cur}\right\|_{2}}$;
9             if $f\left(\Theta\left({\mathbf{x}+r_{best}{\mathbf{d}_{nor}}},\mathbf{x},{\varepsilon_{b}}\right)\right)\neq y$ then
10                   $r_{start},r_{end}\leftarrow 0,\min\left\{r_{best},{\left\|\mathbf{d}_{cur}\right\|_{2}}\right\}$;
11                   while $\left|r_{end}-r_{start}\right|>\gamma$ do
12                         $r_{mid}\leftarrow\left(r_{start}+r_{end}\right)/2$;
13                         if $f\left({\Theta\left({\mathbf{x}+r_{mid}{\mathbf{d}_{nor}}},\mathbf{x},{\varepsilon_{b}}\right)}\right)=y$ then
14                               $r_{start}\leftarrow r_{mid}$;
15                        else
16                               $r_{end}\leftarrow r_{mid}$;
17
18                  if $r_{end}<r_{best}$ then
19                         $r_{best},\mathbf{d}_{best}\leftarrow r_{end},\mathbf{d}_{cur}$;
20
21
22      $s\leftarrow s+1$;
23
$\mathbf{x}_{adv}\leftarrow{\Theta\left({\mathbf{x}+r_{best}\frac{{\mathbf{d_{best}}}}{{{{\left\|{\mathbf{d_{best}}}\right\|}_{2}}}}},\mathbf{x},{\varepsilon_{b}}\right)}$.
Algorithm 1 ISAAC-K.
IV-C ISAAC-N
As a byproduct, we find replacing part-wise non-semantic joints can also bring threats to skeletal models. Therefore, we propose ISAAC-N. For non-key joints obtained in the key joint extraction, we believe that some of these joints are relatively insensitive to the model, i.e., adding perturbations to the non-semantic parts has a relatively small impact on the recognition results. Even from the perspective of human visual perception, the perturbations of non-semantic parts will not affect human subjective judgment, e.g., as for the action of “picking up the mobile phone and taking selfies”, we may pay more attention to whether the arm and head are raised visually than the posture or action of the lower body. If perturbing the joints of the lower body to a large extent can cause misclassification of the model, we deem that the model is unsafe in this context. This is based on the basic fact that a robust model is expected to always focus on the arm joints regardless of the lower body changes i.e., standing, sitting, prostrating.
Taking the NTU120 RGB+D dataset [41] as an example, we find that most of the non-semantic regions are in the lower body, which is also consistent with human cognition since most of the classes in this dataset are upper-body actions. We simply design four typical lower body postures, sit, crouch, kneel, and half kneel, to replace the lower body of the input motion, followed by root joint translation and normalization. If the replaced motion is not adversarial yet, we follow the improved RayS-based perturbation optimization mentioned in Section IV-B. Recall that the temporal consistency constraint helps to improve temporal consistency. This improvement is particularly evident in ISAAC-N when large perturbations are considered. Therefore, we provide three attacking strategies for ISAAC-N: only replacing the lower body (ISAAC-NR), replacement followed by adding perturbations to only the lower body (ISAAC-NRL) and all joints (ISAAC-NRA).
V Experimental Evaluation
In this section, we begin with an experimental setup, followed by a comprehensive evaluation including quantitative and qualitative analyses of ISAAC-K and ISAAC-N, ablation study and countermeasures to show the overall performance of our proposed methods.
V-A Experimental Setup
Datasets. We use three prevalently used datasets, NTU60 RGB+D (NTU60) [36], NTU120 RGB+D (NTU120) [41] and NW-UCLA [42]. NTU60 was captured by three Kinect V2 cameras in 2016 and comprised of 56,880 video clips with 60 classes performed by 40 subjects. The dataset contains 40 everyday actions, 11 two-person interactions, and 9 health actions. Extended by NTU60, NTU120 was released 3 years later and was captured by the same cameras, with another 57,600 video clips and 60 classes, among which 42 classes are everyday actions, 15 classes are two-person interactions, and 3 classes are health actions. The videos were performed by 106 different subjects. NW-UCLA includes 1,494 video clips with 10 action classes. These videos were captured using three Kinect cameras positioned at various viewpoints. Of these views, two are used for training and one for testing. As NTU60 and NTU120 both contain 2D (17 body joints) and 3D (25 body joints) data, we use both of them in our experiments.
Models. We carefully select victim models to show the effectiveness and scalability of our attack. For NTU60 and NTU120, we first select six GCN-based models (MS-AAGCN [43], CTR-GCN [44], DG-STGCN [45], MS-G3D [46], ST-GCN [37]) due to their representativeness. We also involve two SOTA GCN-based models (ST-GCN++ [47], TD-GCN [48], LA-GCN [49]) for NW-UCLA. In recent years, the emergence of Vision Transformer (ViT) has significantly enhanced classification performance in skeleton action recognition tasks. Therefore, we choose two SOTA accessible ViT-based 3D models (Hyperformer [50] and SkateFormer [51]) for all datasets. We train these models from scratch unless the pre-trained models are provided publicly from a public toolbox, PYSKL [52].
Competitors. We compare our method with BASAR [20] since it is the only existing open-source attack under the black-box setting. BASAR uses a random search to find the adversarial motion in the human manifold. For a fair comparison, we modify BASAR from $\ell_{2}$ to $\ell_{\infty}$. We will provide the justification when introducing the metrics later. We do not compare our attack with the later proposed black-box attack, QESAR [21], for the following reasons. 1) QESAR is not open-source, and its reproduction is challenging due to insufficiently clarified constraints in the original paper. 2) QESAR employs a decision-based strategy similar to BASAR, which is time-consuming due to its reliance on random search. As a result, although QESAR claims to reduce the query number compared to BASAR, it still requires an order of 1,000 queries to achieve success, whereas our attack only requires an order of 10 queries (see Section V-B).
Attack Preliminaries. For all datasets, we randomly select 500 skeletal motions respectively. We use ST-GCN as our local surrogate model for extracting key joints. The key joint number is set as 12/18 for 2D/3D data due to better overall performance. We set the restriction threshold $\varepsilon=0.4$ by comprehensively considering the stealthiness-efficiency trade-off. Please refer to Section V-D for more details about different key joint numbers and different $\varepsilon$. Following [18], we set the bone length constraint threshold $\varepsilon_{b}$ as 0.1. The query limit is set as 2,000.
Metrics. We use Attack Success Rate (ASR) and Average Query (AQ) to evaluate the attack performance. ASR denotes the ratio of adversarial skeletons that successfully mislead the classifier. Any attack exceeding the query limit (2,000) will also be considered a failure. AQ stands for the required query number for a successful attack. When ASR is close, lower AQ becomes more significant. We set the restriction threshold $\varepsilon$ (the $\ell_{\infty}$ norm between original and adversarial motion) as 0.4. As a complement to the metrics, following previous works [16, 20], we use the averaged joint acceleration deviation $\Delta a$ (defined as follows) to quantitatively measure the difference between original and adversarial motions.
$\Delta a=\frac{1}{nTN}\sum_{i=1}^{n}\left\|\ddot{\mathbf{x}}_{(i)}-\ddot{\mathbf{x}}^{{}^{\prime}}_{(i)}\right\|_{2},$ (9)
where $\ddot{x}_{(i)}$ denotes the joint acceleration deviation of the $i$-th motion, $N$ denotes the total number of joints in a skeleton, $n$ is the number of adversarial motions. To gauge the distance of joints towards root joint, we also introduce the averaged off-centered deviation $l_{c}$, which can be expressed as
${l_{c}}=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1,j\neq c}^{O}{\left|{{{\left\|{{{\mathbf{x}}_{\left({i,j}\right)}}-{{\mathbf{x}}_{\left({i,c}\right)}}}\right\|}_{2}}-{{\left\|{{\mathbf{x}}_{\left({i,j}\right)}^{{}^{\prime}}-{\mathbf{x}}_{\left({i,c}\right)}^{{}^{\prime}}}\right\|}_{2}}}\right|}},$ (10)
where ${{{\mathbf{x}}_{\left({i,j}\right)}}}$ denotes the $j$-th joint of the $i$-th motion, $c$ is the root joint of the skeleton we select (e.g., joint 21 in the NTU dataset). $l_{c}$ and $\Delta a$ describe the static and dynamic changes of bones, respectively. We replace the averaged joint position deviation $l$ used in BASAR with $l_{c}$ for two reasons. On the one hand, the skeletons may be translated as a whole in the image, resulting in an inaccurate calculation of the absolute distance in $l$. On the other hand, it is unfair to compare the $\ell_{2}$ distance in $l$ since we consider $\ell_{\infty}$ restriction in our attack. Therefore, we calculate the average distance between each joint to the root joint to ensure a fair comparison as much as possible.
TABLE II: Attack performance comparison on NTU60.
Model (Test Accuracy) Attack NTU60(2D) NTU60(3D)
[…]
MS-AAGCN (89.7%/89.0%) BASAR 0.99 51.48 0.08 2.92 0.97 133.58 0.10 3.06
@@ V-A Experimental Setup
Model (Test Accuracy) Attack NTU120(2D) NTU120(3D)
@@
Model (Test Accuracy) Attack NW-UCLA
@@ V-B Attack Performance of ISAAC-K
Attack Results. Table II,  III and  IV report the attack performance of BASAR and ISAAC-K on NTU60, NTU120 and NW-UCLA respectively. The accuracy for the test data is shown under the model name, suggesting that the models are well-trained. In most cases, the ASR of ISAAC-K achieves nearly 100%. However, ISAAC-K consumes much fewer queries than BASAR, indicating that ISAAC-K is more query-efficient. This superiority can be attributed to a smaller search space than boundary-based attack and joint sparsity, where only key joints that have a strong correlation with the output label are perturbed. Although BASAR considers the on-manifold issue, the solution of human body dynamics often falls into a stalemate, resulting in the inability to converge. In terms of quantitative metrics regarding stealthiness, ISAAC-K also performs better than BASAR on $l_{c}$, $\Delta a$ in most cases. Due to the sparsity perturbations, a successful attack may require adding slightly larger perturbations on a limited number of joints, resulting in some cases where the $l_{c}$, $\Delta a$ of ISAAC-K are larger than those of BASAR. However, these figures are still small enough that the perturbations are difficult for human eyes to distinguish, which will be verified at length in Section V-C. Considering that ASR and AQ are the most important indicators to measure attack efficiency, we do not bold these two metrics in tables. Figure 5(a) shows a demonstration of the original and adversarial motions generated by ISAAC-K. The adversarial perturbation is rather small, and the stealthiness can be well-preserved. To verify the universality of this conclusion, Figure 6 provides more visualizations.
@@
Moreover, despite the fact that the key joints of ISAAC-K are extracted from ST-GCN, the competitive attack performance against unseen models indicates that the extracted key joints have a strong transferability across models. Another phenomenon is that attacks against 2D skeletons are found to be easier (fewer queries) than 3D skeletons because of fewer body joints.
Figure 7: Attack performance using different key joints: 1) different key joint numbers, 2) randomly selected key joints.
Figure 8: ASR-Query curves of ISAAC-K and ISAAC-K (all).
Key Joint Analyses. In order to verify that the key joints extracted from Grad-CAM do positively contribute to the attack, we conduct attacks when key joints are selected randomly. Figure 7 shows that, although random selection may be effective, the ASR-query curves converge faster with high ASR as the query increases when the key joints are selected by Grad-CAM. In another aspect, Figure 8 compares the performance of ISAAC-K with a variant, ISAAC-K (all), where all joints are perturbed. It is not surprising that ISAAC-K works better than ISAAC-K (all), illustrating that the key joints extracted from Grad-CAM are indeed those that play crucial roles in classification, thus positively contributing to the attack. The gap between ISAAC-K and ISAAC-K (all) also illustrates that sparse perturbations work better.
@@ V-C User Study
To measure the indistinguishability of adversarial motions, we conducted a user study on Amazon Turk (AMT) [53]. We recruited 80 human subjects to complete three tasks related to adversarial motions: consistency, naturalness, and realness. In our user study, all questions were optional and we did not collect unnecessary personal information. All participants are over 18 years old and consented for their answers to be used for academic research. Our criterion for selecting participants is that their answer approval rate reaches at least 95%, indicating that their previous answers were widely recognized by previous survey initiators. We filter out those arbitrary answers whose answering time is less than eight seconds, since all motions involved in the user study last at least that long. The Human Research Ethics Committee of the authors’ affiliation determined that the study was exempt from further human subjects review, and we followed best practices for ethical human subjects survey research.
@@
Survey Protocol. 1) Consistency. Whether the skeletal motion is consistent to the output label plays a dominant role. We randomly select 10 clean motions and 60 adversarial motions averagely from six GCN-based victim models (10 from each) and provide their original labels and adversarial labels. Subjects are asked which label best describes the skeletal motion. 2) Naturalness. Although we set a restriction threshold to the adversarial motion, whether the perturbations are stealthy enough to maintain the visual naturalness is still hard to measure quantitatively. Instead, following prior work [27], the subjects are asked to use a Likert-scale [54] from 1 to 5 (higher means more natural) to evaluate the naturalness of randomly selected 60 skeletal motions which consists of 30 clean ones, and 30 adversarial ones. 3) Realness. It is also a challenging task to distinguish which skeletal motion is real (clean) if two motions are played together. The subjects are provided with two motions (a clean one and an adversarial one) with minor differences and asked to choose which skeletal motion is the real one.
Figure 9: Naturalness test results.
@@
Results and Analyses. In the consistency test, we find almost all the subjects (92.5%) could not choose the original label for adversarial motions, indicating that the perturbations are small enough to preserve the semantic information of the motions. Figure 9 shows the naturalness test results. The subjects gave similar results between clean and adversarial motions. We conduct a Mann-Whitney U-test [55] between the two distributions, with the null hypothesis being that there is no significant difference between original and adversarial motions in terms of naturalness. The result (p = 0.08$>$0.05) indicates that ISAAC-K maintains sensory comfort well. Since we modify BASAR to apply to more practical scenarios, the naturalness drops a lot. In the realness test, we find that, overall, 32.19% of the subjects could correctly distinguish adversarial motions from clean ones, which is lower than random guessing. It can be concluded that ISAAC-K can not only maintain semantic information similar to the original motions, but also confuse human subjects in distinguishability.
@@ V-D Ablation Study
Restriction Threshold $\varepsilon$. Changing restriction threshold $\varepsilon$ will affect the attack performance as well as, more importantly, the naturalness of the adversarial motions. Figure 10 shows the ASR-query curves under different restriction thresholds $\varepsilon$. ASR increases sharply and then slows down as $\varepsilon$ increases. After $\varepsilon$ exceeds 0.4, ASR increases limitedly and the artifacts become obvious in adversarial motions (the average naturalness drops below “neutral” in Figure 11), providing a reasonable explanation of why we choose $\varepsilon$ as 0.4 finally.
Figure 10: Attack performance test with different $\varepsilon$.
Figure 11: Naturalness test with different $\varepsilon$.
Key Joint Number $N_{k}$. We also conduct an ablation study about different key joint numbers $N_{k}$. Figure 7 shows the ASR-query curves. The results are consistent with our intuition that a small key joint number will increase the attack difficulty (lower ASR), and a larger key joint number will reduce the attack efficiency (higher AQ). Users can choose different $N_{k}$ according to their preferences or requirements, given insignificant differences.
@@ V-E Attack Performance of ISAAC-N
Figure 5(b) visualizes ISAAC-NR by an actual example. The replacement does not change the semantic information of the motion and maintains the naturalness of the adversarial motion. The naturalness test result in Figure 9 shows that ISAAC-N obtains sound insidiousness (with a p-value of 0.06$>$0.05) since the perturbations are generated by solely replacing the lower body. Figure 12 provides more visualizations for ISAAC-N.
TABLE V: Attack performance comparison for ISAAC-N on NTU120(3D).
@@ V-F Countermeasure
Performance against the Existing Defense. To test the ability of our attack to resist defense, we employ a SOTA defense, BEAT [31], which is a Bayesian-based AT method. We do not validate our attack on 2D data, since BEAT is tailored for 3D data. Table VI reports the performance of re-attacking the trained defense models by ISAAC-K. The ASR and AQ are close to attacking the original target model before adversarial training, indicating that BEAT is almost unable to defend against ISAAC-K. Considering the performance of BASAR under BEAT (verified in the paper of BEAT [31]), we find that BEAT can alleviate BASAR, but BEAT’s ability to defend against ISAAC-K is very limited. One possible reason could be that BASAR and BEAT are a combination of attack and defense that focuses on perturbation stealthiness rather than attack efficiency. In addition, we do not consider the forgery detection [33] since it primarily targets additional frames and actions inserted into a sequence of skeleton actions. Therefore, it does not fall within the realm of adversarial attacks. We also dismiss the Fourier-based AT [34] since it is designed only for $\ell_{2}$-PGD attacks (our attack uses $\ell_{\infty}$) and is not open-source.
TABLE VI: Performance of ISAAC-K before/after BEAT.
@@
Adaptive Defense. We also show that our proposed attack can in return help improve the adversarial robustness. As AT has been universally acknowledged as the most feasible defense in theory, we ameliorate AT and propose several adaptive defenses based on adversarial motions generated by ISAAC-K and ISAAC-N. Generally, given dataset $\mathcal{D}$, we solve the following min-max problem for model $f$ with parameters $\theta$.
$\mathop{\min}\limits_{\theta}{\mathop{\mathbbm{missing}}{E}\limits_{\left({{\mathbf{x}},y}\right)\sim\mathcal{D}}}\left[{\mathop{\max}\limits_{r,{\mathbf{d}}}L\left[{\Theta\left({{\varphi\left({\mathbf{x}}\right)}+r\delta\frac{{{M_{x}}\odot{\mathbf{d}}}}{{{{\left\|{{M_{x}}\odot{\mathbf{d}}}\right\|}_{2}}}},{\varphi\left({\mathbf{x}}\right)},{\varepsilon_{b}}}\right),y}\right]}\right].$ (11)
@@
    AT-ISAAC-NRL: $\varphi\left({\mathbf{x}}\right)={\mathcal{R}}\left({{\mathbf{x}},{{\mathbf{x}}_{r}},{M_{r}}}\right)$, $\delta=1$, ${M_{x}}={M_{r}},\\ \mathop{\max}\limits_{r,{\mathbf{d}}}\to{\mathbbm{E}_{\left({{{\mathbf{x}}_{r}},{M_{r}}}\right)}}\mathop{\max}\limits_{r,{\mathbf{d}}}$.
@@
    AT-ISAAC-NRA: $\varphi\left({\mathbf{x}}\right)={\mathcal{R}}\left({{\mathbf{x}},{{\mathbf{x}}_{r}},{M_{r}}}\right)$, $\delta=1$, ${M_{x}}=\mathbf{1}_{T\times N\times D},\\ \mathop{\max}\limits_{r,{\mathbf{d}}}\to{\mathbbm{E}_{\left({{{\mathbf{x}}_{r}},{M_{r}}}\right)}}\mathop{\max}\limits_{r,{\mathbf{d}}}$.
@@
We train four ST-GCN++ models with AT, and we only consider the lower body of ‘sit’ in the last three versions. The results in Table VII demonstrate a certain defense capability of these four adaptive defenses. After AT, ASR drops in all defenses, among which AT-ISAAC-NRA contributes the most to ASR reduction. AQ also increases after AT, indicating the increasing difficulty for re-attacks.

### v2

@@ V-A Experimental Setup
MS-AAGCN
(89.7%/89.0%)
BASAR 0.99 51.48 0.08 2.92 0.97 133.58 0.10 3.06
@@
Model
(Test Accuracy)
Attack NTU120(2D) NTU120(3D)
@@
Model
(Test Accuracy)
Attack NW-UCLA
@@ V-B Attack Performance of ISAAC-K
Attack Results. Table IIIII and IV report the attack performance of BASAR and ISAAC-K on NTU60, NTU120 and NW-UCLA respectively. The accuracy for the test data is shown under the model name, suggesting that the models are well-trained. In most cases, the ASR of ISAAC-K achieves nearly 100%. However, ISAAC-K consumes much fewer queries than BASAR, indicating that ISAAC-K is more query-efficient. This superiority can be attributed to a smaller search space than boundary-based attack and joint sparsity, where only key joints that have a strong correlation with the output label are perturbed. Although BASAR considers the on-manifold issue, the solution of human body dynamics often falls into a stalemate, resulting in the inability to converge. In terms of quantitative metrics regarding stealthiness, ISAAC-K also performs better than BASAR on $l_{c}$ and $\Delta a$ in most cases. Due to the sparsity of perturbations, a successful attack may require the addition of slightly larger perturbations to a limited number of joints, resulting in some cases where the metrics are larger than those of BASAR. However, these figures are still small enough that the perturbations are difficult for human eyes to distinguish, which will be verified at length in Section V-C. Considering that ASR and AQ are the most important indicators to measure attack efficiency, we do not bold these two metrics in tables. Fig. 5(a) shows an example of the original and adversarial motions generated by ISAAC-K. The adversarial perturbation is rather small, and the stealthiness can be well-preserved. To verify the universality of this conclusion, Fig. 6 provides more visualizations.
@@
Moreover, although the key joints of ISAAC-K are extracted from ST-GCN, the competitive attack performance against unseen models indicates that the extracted key joints have a strong transferability across models. Another phenomenon is that attacks against 2D skeletons are found to be easier (fewer queries) than 3D skeletons because of fewer body joints.
Figure 7: Attack performance under key joint selection and random joint selection across different joint numbers.
Figure 8: Attack performance of ISAAC-K and ISAAC-K (all).
Key Joint Analyses. In order to verify that the key joints extracted from Grad-CAM do positively contribute to the attack, we conduct attacks when key joints are selected randomly. Fig. 7 shows that, although random selection may be effective, the ASR-query curves converge faster with high ASR as the query increases when the key joints are selected by Grad-CAM. In another aspect, Fig. 8 compares the performance of ISAAC-K with a variant, ISAAC-K (all), where all joints are perturbed. It is not surprising that ISAAC-K performs better than ISAAC-K (all), illustrating that the key joints extracted from Grad-CAM are indeed those that play crucial roles in classification, thus positively contributing to the attack. The gap between ISAAC-K and ISAAC-K (all) also illustrates that sparse perturbations work better.
@@ V-C User Study
To measure the indistinguishability of adversarial motions, we conducted a user study on Amazon Turk (AMT) [amazon_turk]. We recruited 80 human subjects to complete three tasks related to adversarial motions: consistency, naturalness, and realness. In our user study, all questions were optional and we did not collect unnecessary personal information. All participants were over 18 years old and consented for their answers to be used for academic research. Our criterion for selecting participants is that their answer approval rate reaches at least 95%, indicating that their previous answers were widely recognized by previous survey initiators. We filter out those arbitrary answers whose answering time is less than eight seconds, since all motions involved in the user study last at least that long. The Human Research Ethics Committee of the authors’ affiliation determined that the study was exempt from further human subjects review, and we followed best practices for ethical human subjects survey research.
@@
Survey Protocol. 1) Consistency. Whether the skeletal motion is consistent to the output label plays a dominant role. We randomly selected 10 clean motions and 60 adversarial motions averagely from six GCN-based victim models (10 from each) and provided their original labels and adversarial labels. Subjects were asked to select the label that best describes the skeletal motion. 2) Naturalness. Although we set a restriction threshold on the adversarial motion, it remains difficult to quantitatively assess whether the perturbations are stealthy enough to preserve visual naturalness. Instead, following prior work [cao2023stylefool], the subjects were asked to use a Likert-scale [likert1932technique] from 1 to 5 (higher means more natural) to evaluate the naturalness of randomly selected 60 skeletal motions which consists of 30 clean ones, and 30 adversarial ones. 3) Realness. It is also a challenging task to distinguish which skeletal motion is real (clean) if two motions are played together. The subjects were provided with two motions (a clean one and an adversarial one) with minor differences and asked to choose which skeletal motion is the real one.
Figure 9: Naturalness test results of different attacks.
@@
Results and Analyses. In the consistency test, we find almost all the subjects (92.5%) chose the original label for adversarial motions, indicating that the perturbations are small enough to preserve the semantic information of the motions. Fig. 9 shows the naturalness test results. The subjects gave similar results between clean and adversarial motions. We conduct a Mann-Whitney U-test [mann1947on] between the two distributions, with the null hypothesis being that there is no significant difference between original and adversarial motions in terms of naturalness. The result (p = 0.08>>0.05) indicates that ISAAC-K maintains sensory comfort well. Since we modify BASAR to apply to more practical scenarios, the naturalness drops a lot. In the realness test, we find that, overall, 32.19% of the subjects could correctly distinguish adversarial motions from clean ones, which is lower than random guessing. It can be concluded that ISAAC-K can not only maintain semantic information similar to the original motions, but also confuse human subjects visually.
@@ V-D Ablation Study
Restriction Threshold $\varepsilon$. Changing the restriction threshold $\varepsilon$ will affect the attack performance as well as, more importantly, the naturalness of the adversarial motions. Fig. 10 shows the ASR-query curves under different restriction thresholds $\varepsilon$. ASR increases sharply and then slows down as $\varepsilon$ increases. After $\varepsilon$ exceeds 0.4, ASR increases limitedly and the artifacts become obvious in adversarial motions (the average naturalness drops below “neutral” in Fig. 11), providing a reasonable explanation of why we choose $\varepsilon$ as 0.4 finally.
Figure 10: Attack performance under different restriction thresholds $\varepsilon$.
Figure 11: Naturalness test results under different restriction thresholds $\varepsilon$.
Key Joint Number $N_{k}$. We also conduct an ablation study under different key joint numbers $N_{k}$. Fig. 7 shows the ASR-query curves. The results are consistent with our intuition that a small key joint number will increase the attack difficulty (lower ASR), and a larger key joint number will reduce the attack efficiency (higher AQ). Users can choose different $N_{k}$ according to their preferences or requirements, given insignificant differences.
@@ V-E Attack Performance of ISAAC-N
Fig. 5(b) visualizes ISAAC-NR by an actual example. The replacement preserves the semantic information of the motion and maintains the naturalness of the adversarial motion. The naturalness test result in Fig. 9 shows that ISAAC-N obtains sound insidiousness (with a p-value of 0.06>>0.05) since the perturbations are generated by solely replacing the lower body. Fig. 12 provides more visualizations for ISAAC-N.
TABLE V: Attack performance of ISAAC-N on NTU120(3D).
@@ V-F Countermeasure
Performance against the Existing Defense. To test the ability of our attack to resist defense, we employ a SOTA skeletal defense, BEAT [wang2023defending], which is a Bayesian-based AT method. We do not validate our attack on 2D data, since BEAT is tailored for 3D data. Table VI reports the performance of re-attacking the trained defense models by ISAAC-K. Due to the differences in perturbation constraints and sparsity levels between the adversarial examples used in BEAT and those generated by ISAAC-K, the ASR and AQ are close to those before adversarial training, indicating that BEAT is almost unable to defend against ISAAC-K. Considering the performance of BASAR under BEAT (verified in the paper of BEAT [wang2023defending]), we find that BEAT can alleviate BASAR, but BEAT’s ability to defend against ISAAC-K is very limited. One possible reason could be that BASAR and BEAT are a combination of attack and defense that focuses on perturbation stealthiness rather than attack efficiency.
TABLE VI: Performance of ISAAC-K against BEAT and RS.
@@
Furthermore, we evaluate the robustness of our attacks against a defense based on randomized smoothing (RS) [zheng2020towards]. We extend the RS framework from image classification to skeleton action recognition by incorporating Gaussian noise and temporal filtering. This method computes a certified robustness radius based on the variance of perturbed samples, and considers an adversarial skeleton to be defended if its distance from the clean skeleton lies within this radius. As shown in Table VI, the Defense Success Rate (DSR) of RS is consistently close to 0 against our attacks. This can be attributed to two factors: (1) our perturbations are crafted under an $\ell_{\infty}$-norm constraint, for which RS does not provide certification guarantees [kumar2020curse]; and (2) skeleton data is inherently sparse, topologically discrete, and low-dimensional. Therefore, skeleton models are less sensitive to small perturbations than image models. These characteristics limit the effectiveness of RS-based certification on skeleton models.
Adaptive Defense. We also show that our proposed attack can in return help improve the adversarial robustness. As AT has been universally acknowledged as the most feasible defense in theory, we ameliorate AT and propose several adaptive defenses based on adversarial motions generated by ISAAC-K and ISAAC-N. Generally, given dataset $\mathcal{D}$ and cross-entropy loss function $\widetilde{\ell}$, we solve the following min-max problem for model $f$ with parameters $\theta$.
$\small\mathop{\min}\limits_{\theta}{\mathop{\mathbbm{E}}\limits_{\left({{\mathbf{x}},y}\right)\sim\mathcal{D}}}\left[{\mathop{\max}\limits_{r,{\mathbf{d}}}\widetilde{\ell}\left[{\Theta\left({{\varphi\left({\mathbf{x}}\right)}+r\delta\frac{{{M_{x}}\odot{\mathbf{d}}}}{{{{\left\|{{M_{x}}\odot{\mathbf{d}}}\right\|}_{2}}}},{\varphi\left({\mathbf{x}}\right)},{\varepsilon_{b}}}\right),y}\right]}\right].$ (13)
@@
    AT-ISAAC-NRL: $\varphi\left({\mathbf{x}}\right)={\mathcal{R}}\left({{\mathbf{x}},{{\mathbf{x}}_{r}},{M_{r}}}\right)$, $\delta=1$, ${M_{x}}={M_{r}},\mathop{\max}\limits_{r,{\mathbf{d}}}\to{\mathbbm{E}_{\left({{{\mathbf{x}}_{r}},{M_{r}}}\right)}}\mathop{\max}\limits_{r,{\mathbf{d}}}$.
@@
    AT-ISAAC-NRA: $\varphi\left({\mathbf{x}}\right)={\mathcal{R}}\left({{\mathbf{x}},{{\mathbf{x}}_{r}},{M_{r}}}\right)$, $\delta=1$, ${M_{x}}=\mathbf{1}_{T\times N\times D},\mathop{\max}\limits_{r,{\mathbf{d}}}\to{\mathbbm{E}_{\left({{{\mathbf{x}}_{r}},{M_{r}}}\right)}}\mathop{\max}\limits_{r,{\mathbf{d}}}$.
To characterize how our designed defenses influence model generalization under structured adversarial perturbations, we derive a unified robust generalization bound based on the Rademacher complexity [yin2019rademacher]. Given a hypothesis class $\mathcal{F}$, for any $\eta\in(0,1)$, with probability at least $1-\eta$, the following inequality holds for all $f\in\mathcal{F}$,
$\small\widetilde{R}(f)\leq\widetilde{R}_{n}(f)+2\mathfrak{R}_{\mathcal{S}}(\widetilde{\ell}_{\mathcal{F}})+3\sqrt{\frac{\log(2/\eta)}{2|\mathcal{D}|}},$ (14)
[…]
@@
We train four ST-GCN++ models with AT, and we only consider the lower body of “sit” in the last three versions. During training, we use examples optimized by 5 PGD steps [madry2018towards]. The results in Table VII demonstrate a certain defense capability of these four adaptive defenses. After AT, ASR drops in all defenses, among which AT-ISAAC-NRA contributes the most to ASR reduction. AQ also increases after AT, indicating the increasing difficulty for re-attacks. However, we find that AT still remains insufficient for our attacks. One possible reason is that, unlike image models where 5 PGD steps usually push samples close to the decision boundary, our attacks on sparse and low-dimensional skeleton data with additional kinematic and consistency constraints require much deeper optimization to make the skeleton truly adversarial. Consequently, the PGD samples used during training often remain far from the decision boundary, and the mismatch between their distribution and that of our attacks further limits the effectiveness of AT.